# System Identification for Dynamic Modeling of a Bumper Car

### v1

@@ 5 Conclusion
The authors would like to thank Yannick Soller for his technical assistance with the experimental setup. His support in collecting data on the real system is gratefully acknowledged.
Appendix A State Estimation
Appendix B Genetic algorithm for parameter optimization of the Newton-Euler model
@@
Maximum stepper speed $d$ 1.418 rad/sec
@@
Throttle proportional gain* $K_{\mathrm{t}}$ 274.8 Nsec/m
@@
Appendix C Hyperparameters of MLP-training
Appendix D SINDy regression and model parameters
We use the SINDy framework implemented in PySINDy (de Silva et al., 2020; Kaptanoglu et al., 2022). All states entering the candidate matrices are scaled accordingly to ensure consistent thresholding.
@@ Appendix D SINDy regression and model parameters
Appendix E Discussion on predictive accuracy

### v2

@@ 5 Conclusion
The authors would like to thank Yannick Soller for his technical assistance with the experimental setup and support in collecting data on the real system.
Appendix A Actuator dynamics
with maximum steer rate $d>0$ and time constant $T_{\mathrm{s}}>0$.
Appendix B Tire Slip computation
For tire slip computation, Kim et al. (2018) introduced an Advanced Slip Ratio (ASR), where a lower bound is imposed on the longitudinal velocity when computing the tire slip angle, thus implying $\alpha\to 0$ for low velocities. Moreover, the extended steering range of a bumper cars induces negative longitudinal velocity, causing the rear wheel to rotate in reverse. As a result, the rear tire slip is given by $\alpha_{\mathrm{r}}=\beta_{\mathrm{r}}-\delta_{\mathrm{r}}$ where $\delta_{\mathrm{r}}$ is 0° if the wheel spins forward and 180° if backward. Due to the trigonometric properties of the $\arctan(\cdot)$ function, we compute the tire heading angle using the ASR and the absolute value of the longitudinal tire velocity as
Kim et al. (2018) shows that the integration of ASR limits the lateral tire dynamics, however, the lateral tire dynamics remain faster than the dynamics of the vehicle body inertia, making the Newton-Euler model stiff.
Appendix C Hyperparameters of MLP-training
Appendix D Genetic algorithm for parameter optimization of the Newton-Euler model
@@
Maximum stepper speed $d$ 1.418 rad/s
@@
Throttle proportional gain* $K_{\mathrm{t}}$ 274.8 Ns/m
@@
Appendix E SINDy regression and model parameters
We use the SINDy framework implemented in PySINDy (de Silva et al., 2020). All states entering the candidate matrices are scaled accordingly to ensure consistent thresholding.
@@ Appendix E SINDy regression and model parameters
Appendix F State Estimation
Appendix G Discussion on predictive accuracy